OpenStudy (anonymous):

Find the slope of the line that passes through (3, -4) and (-1, 2). 2/3 3/2 -2/3 -3/2

OpenStudy (anonymous):

\[ m= \frac{ Y _{2} -Y _{1}}{ X _{2}-X _{1} }\]

OpenStudy (anonymous):

\[m = \frac{ y2 - y1 }{ x2 - x1 }\] or \[m = \frac{ 2 - (-4) }{ -1 - 3 }\]